# Supernova-induced binary-interaction-powered supernovae: a model for SN2022jli

### v1

@@ Abstract
We present 3D hydrodynamical modelling of supernova-induced binary-interaction-powered supernovae; a scenario proposed for the peculiar type Ic supernova SN2022jli. In this scenario, supernova ejecta of a stripped-envelope star impacts a close-by stellar companion, temporarily inflating the envelope. The expanded envelope engulfs the neutron star, causing strong mass accretion at super-Eddington rates. Feedback from the accretion powers the supernova light curve with periodic undulations. Our simulations capture key features of SN2022jli, both the overall decline and the superimposed undulations of the light curve. Based on our parameter study, we find that (i) the accretion feedback should be sufficiently geometrically confined and the eccentricity of the post-supernova binary orbit should be $0.7\lesssim e\lesssim 0.9$ to sustain a high accretion rate, and (ii) the viewing angle should be close to edge-on to match the low undulation amplitude ($\Delta L/L\sim 0.1$) of SN2022jli. Different combinations of parameters could account for other supernovae like SN2022mop, SN2009ip and SN2015ap. We also discuss possible explanations for other key features of SN2022jli such as the $\gamma$-ray detection at $\sim 200~{}\mathrm{d}$ and the rapid optical drop at $\sim 250~{}\mathrm{d}$. Finally, we speculate on the future evolution of the system and its relation to existing neutron star binaries.
@@ 1 Introduction
The colour in the second portion of the light curve was much bluer than other known stripped-envelope SNe, indicating an extra powering mechanism (Cartier et al., 2024). Along with the clear periodicity, it was speculated that this SN is binary-interaction-powered as proposed in Hirai & Podsiadlowski (2022). In this scenario, the nascent neutron star (NS) directly penetrates the outer layers of a main sequence companion star, causing strong mass accretion that powers the light curve from inside the ejecta. However, such direct interactions can only occur several times as the orbit will rapidly decay upon each penetration, whereas SN2022jli showed an almost fixed periodicity for at least $\gtrsim 15$ cycles.
In this paper, we model a scenario we proposed in previous studies (Ogata et al., 2021; Moore et al., 2023; Hirai, 2024) with 3D hydrodynamic simulations. It builds on the scenario of Hirai & Podsiadlowski (2022) but with the additional effect of ejecta-companion interaction, where the outer layers of the companion star is heated by the impact of the SN ejecta and forced to inflate (e.g. Hirai et al., 2018). With sufficient SN-heating, the companion can inflate enough to overflow its Roche lobe or engulf the entire post-SN orbit, inducing binary interactions. The density of the inflated layers are typically low enough to avoid any measurable amount of orbital decay, but simultaneously high enough to enable strong mass accretion onto the remnant compact object. Since the binary orbit is almost guaranteed to be eccentric due to the SN mass loss and natal kick, we expect there to be periodic modulations in the accretion rate as the accretor travels through various depths of the companion envelope. The envelope is sufficiently optically thick to reprocess any radiative and kinetic feedback from the accretor into optical wavelengths. For all of the above reasons, this is an ideal situation to create bright undulating light curves like that of SN2022jli.
This paper is organized as follows. We first outline our model parameters and the methods for the hydrodynamic simulations and light curve modelling in Section 2. Simulation results are presented Section 3, along with a comparison of the simulated light curve with the observed light curve of SN2022jli. We then discuss how to explain other key features of SN2022jli within our scenario as well as speculations on the future evolution of the system in Section 4. We finally summarize our results in Section 5.
@@ 2.1 Binary parameters
There are several possible constraints on the companion mass. Since the SN was classified as type Ic and the SN2022jli progenitor system likely had a pre-SN orbital period of $P_{\mathrm{orb}}<12.5~{}\mathrm{d}$, it most likely experienced a common-envelope phase that stripped the primary envelope and tightened the orbit. Although the exact conditions for initiating common-envelope phases are not known (e.g. Ivanova et al., 2013; Pavlovskii et al., 2017; Temmink et al., 2023; Willcox et al., 2023), it is considered that low mass ratio systems are generally more unstable. This places a loose upper bound on the companion mass to be sufficiently smaller than the pre-common-envelope primary mass. On the other hand, the companion mass should be large enough such that the binary can stay bound after the SN. In the absence of NS natal kicks, the binary will survive when the ejecta mass is less than half of the total mass of the binary. The estimated ejecta mass for SN2022jli is $M_{\mathrm{ej}}\sim 1.5~{}M_{\odot}$ (Chen et al., 2024; Cartier et al., 2024), so assuming a NS mass of $M_{\mathrm{NS}}=1.4~{}M_{\odot}$, we cannot place a useful lower limit on the companion mass. Given these considerations, we choose a companion mass of $M_{2}=5~{}M_{\odot}$ as our fiducial model. For comparison, we also run models with companion masses of $M_{2}=3~{}M_{\odot}$ and $10~{}M_{\odot}$. All stellar models are constructed with the stellar evolution code MESA (Paxton et al., 2011), assuming that the companion is close to its zero-age main-sequence structure.
For the orbital eccentricity, we choose $e=0.5$ as our baseline model and run $e=0.3$ and $0.7$ for comparison. All model parameters are summarized in Table 1. If we assume the pre-SN orbit was circular and that the ejecta mass was $M_{\mathrm{ej}}=1.5~{}M_{\odot}$, the mass loss alone can only excite an eccentricity of $e=M_{\mathrm{ej}}/(M_{\mathrm{NS}}+M_{2})=1/3,3/13,2/23$ for $M_{2}=3,5,10~{}M_{\odot}$ respectively. Our choices of eccentricity implies that there was also a natal kick imparted to the NS with a magnitude of the order of $v_{\mathrm{kick}}\gtrsim 50~{}\mathrm{km~{}s}^{-1}$.
@@
Name $M_{2}$ $e$ $a_{\mathrm{per}}$ Feedback
@@
M05e05b 5 0.5 21.1 Bipolar
M05e03b 5 0.3 29.6 Bipolar
M05e07b 5 0.7 12.7 Bipolar
M10e05b 10 0.5 25.6 Bipolar
M03e05b 3 0.5 18.7 Bipolar
M05e05n 5 0.5 21.1 None
M05e05t 5 0.5 21.1 Thermal
@@ 2.2 Hydrodynamics code
We use a spherical coordinate system for all of our simulations. The companion star model is centered on the origin. Since our code cannot handle vacuum, we attach a dilute atmosphere outside the star, with a density profile $\rho=\rho_{0}(R_{2}/r)^{2}$, where $\rho_{0}$ is set to a value lower than the surface density of the companion. The outer boundary extends out to $r_{\mathrm{out}}=8\times 10^{15}~{}\mathrm{cm}\sim 115,000~{}R_{\odot}$, which is $\sim 3000$ times larger than the orbital semimajor axis. The radial grid spacing is chosen such that it is smallest around the surface of the star and increases in a geometric series inwards and outwards. This ensures that the pressure scale height in the star is always resolved by at least $\sim 10$ grid points [^3] Except for the very surface layers where the pressure scale height is prohibitively small.. We cover the star with $\sim 300$ radial grid points, and $N_{r}=800$ grid points altogether. We assume equatorial symmetry, and thus only simulate the upper hemisphere where the polar grid spacing is uniform in $\cos\theta$, with $N_{\theta}=32$ grid points from pole to equator. This ensures that the solid angle per cell is uniform, minimizing asymmetries in numerical viscosity. It also places more resolution around the orbital plane, where most of the interesting dynamics are expected to take place. The azimuthal spacing is uniform with $N_{\phi}=256$ grid points. To relax the severe Courant conditions around the coordinate origin, we gradually reduce the angular resolution towards the centre, assuming full spherical symmetry ($N_{\theta}=N_{\phi}=1$) in the innermost 5 cells, $(N_{\theta},N_{\phi})=(2,16)$ in the next 5 cells, $(N_{\theta},N_{\phi})=(4,32)$ in the next 6 cells and so on.
@@ 2.3 SN-heating
After the SN explosion, the SN ejecta collide with the companion star, heating the surface layers through shocks. The injected heat drives the surface out of thermal equilibrium, causing it to become overluminous and inflated (Hirai & Yamada, 2015; Hirai et al., 2018; Ogata et al., 2021; Chen et al., 2023; Hirai, 2023). Depending on the amount of energy injection and the stellar mass, the star can reach radii of a few ×$\times 100~{}R_{\odot}$. This is far beyond the orbital semi-major axis, inevitably causing direct interactions between the NS and the inflated envelope.
@@
[…]
@@ 2.4 Feedback method
The accretion and feedback is treated in our code as follows. For all cells within a specified accretion radius $|\bm{r}-\bm{r}_{\mathrm{NS}}|\leq R_{\mathrm{acc}}$, we add a source term to the continuity equation in the form
$\frac{\partial\ln\rho}{\partial t}=-\frac{1}{t_{\mathrm{ff}}},$ (4)
where $\rho$ is the density and $t_{\mathrm{ff}}$ is the free-fall time defined by
@@
where $\phi_{\mathrm{soft}}(r)$ is the softened gravitational potential of the NS. Eq. (4) is based on the assumption that the mass inside the accretion radius will be accreted by the NS on the local free-fall timescale. The mass and momentum lost in the process is added to the NS particle to represent accretion. We then update the internal energy density of each cell such that the specific entropy is unchanged after applying the density source term.
For the feedback models, we further calculate the accretion luminosity as
@@
where $G$ is the gravitational constant, $\dot{M}_{\mathrm{NS}}$ is the mass accretion rate and we choose the NS radius to be $R_{\mathrm{NS}}=12~{}\mathrm{km}$. In the thermal feedback model, we assume that all of the accretion energy is liberated in the form of radiation, and a fraction of it is captured by the gas and thermalizes. We define the optical depth within the accretion radius as
@@ 2.5 Light curve modelling
We use the post-processing module developed for HORMONE in Grishin et al. (2021) to compute synthetic light curves. It follows the method of Suzuki et al. (2019) where we shoot a grid of rays from behind the source towards the observer and solve the radiative transfer equations to obtain the intensity along each ray. The intensity is integrated over the whole projected area of the source and multiplied by $4\pi$ to obtain the isotropic-equivalent luminosity $L$. We separately carry out the same shooting procedure for specific intensities at 6 different frequencies to get a rough spectrum. We fit the spectrum with a black body to compute an effective temperature $T_{\mathrm{eff}}$, which is then used to define a photospheric radius $R_{\mathrm{eff}}=\sqrt{L/(4\pi\sigma T_{\mathrm{eff}}^{4})}$ where $\sigma$ is the Stefan-Boltzmann constant.
@@
Given the high asymmetry of our problem, we also investigate the viewing direction dependence. Defining a Cartesian coordinate system with the orbital angular momentum vector as the $z$ axis, the direction opposite of the orbital eccentricity vector (Laplace-Runge-Lenz vector) as the $x$ axis, and the direction of the initial NS velocity as the $y$ axis, we choose 5 different viewing directions; $+z,+x,-x,+y,-y$ (see the arrows in Figure 1). Note that the $+z$ and $-z$ directions are identical as we assume equatorial symmetry. For the ±$\pm x,\pm y$ directions, we compute the light curve from two different polar angles, $\theta=\pi/2$ (edge on) and $\pi/4$ (off plane).
We do not simulate the SN explosion itself, which would be contributing significantly to the light curve, especially at early times. We discuss implications of this in Section 4.2.

### v2

@@ Abstract
We present 3D hydrodynamical modelling of supernova-induced binary-interaction-powered supernovae; a scenario proposed for the peculiar type Ic supernova SN2022jli. In this scenario, supernova ejecta of a stripped-envelope star impact a close-by stellar companion, temporarily inflating the envelope. The expanded envelope engulfs the neutron star, causing strong mass accretion at super-Eddington rates. Feedback from the accretion powers the supernova light curve with periodic undulations. Our simulations capture key features of SN2022jli, both the overall decline and the superimposed undulations of the light curve. Based on our parameter study, we find that (i) the accretion feedback should be sufficiently geometrically confined and (ii) the eccentricity of the post-supernova binary orbit should be $0.8\lesssim e\lesssim 0.9$ to sustain a high accretion rate and match the low undulation amplitude ($\Delta L/L\sim 0.1$) of SN2022jli. Different combinations of parameters could account for other supernovae like SN2022mop, SN2009ip and SN2015ap, which have varying undulation periods and amplitudes. We also discuss possible explanations for other key features of SN2022jli such as the $\gamma$-ray detection at $\sim 200~\mathrm{d}$ and the rapid optical drop at $\sim 250~\mathrm{d}$. Finally, we speculate on the future evolution of the system and its relation to existing neutron star binaries.
@@ 1 Introduction
The colour in the second portion of the light curve was much bluer than other known stripped-envelope SNe, indicating an extra powering mechanism (Cartier et al., 2024). Some studies propose magnetar spin-down as the source of the extra power (Cartier et al., 2024; Orellana et al., 2025), although it requires an additional mechanism to explain the undulations as it does not naturally have any periodicity. Due to the clear periodicity, it was speculated that this SN is binary-interaction-powered as proposed in Hirai & Podsiadlowski (2022). In this scenario, the nascent neutron star (NS) directly penetrates the outer layers of a main sequence companion star, causing strong mass accretion that powers the light curve from inside the ejecta. However, such direct interactions can only occur several times as the orbit will rapidly decay upon each penetration, whereas SN2022jli showed an almost fixed periodicity for at least $\gtrsim 15$ cycles.
In this paper, we model a scenario we proposed in previous studies (Ogata et al., 2021; Moore et al., 2023; Hirai, 2024) with 3D hydrodynamic simulations. It builds on the scenario of Hirai & Podsiadlowski (2022) but with the additional effect of ejecta-companion interaction, where the outer layers of the companion star is heated by the impact of the SN ejecta and forced to inflate (e.g. Hirai et al., 2018). With sufficient SN-heating, the companion can inflate enough to overflow its Roche lobe or engulf the entire post-SN orbit, inducing binary interactions. The density of the inflated layers are typically low enough to avoid any measurable amount of orbital decay, but simultaneously high enough to enable strong mass accretion onto the remnant compact object (Hober et al., 2022; Soker, 2022). Since the binary orbit is almost guaranteed to be eccentric due to the SN mass loss and natal kick, we expect there to be periodic modulations in the accretion rate as the accretor travels through various depths of the companion envelope. The envelope is sufficiently optically thick to reprocess any radiative and kinetic feedback from the accretor into optical wavelengths. For all of the above reasons, this is an ideal situation to create bright undulating light curves like that of SN2022jli.
This paper is organized as follows. We first outline our model parameters and the methods for the hydrodynamic simulations and light curve modelling in Section 2. Simulation results are presented in Section 3, along with a comparison of the simulated light curve with the observed light curve of SN2022jli. We then discuss how to explain other key features of SN2022jli within our scenario as well as speculations on the future evolution of the system in Section 4. We finally summarize our results in Section 5.
@@ 2.1 Binary parameters
There are several possible constraints on the companion mass. Since the SN was classified as type Ic and the SN2022jli progenitor system likely had a pre-SN orbital period of $P_{\mathrm{orb}}<12.5~\mathrm{d}$, it most likely experienced a common-envelope phase that stripped the primary envelope and tightened the orbit. Although the exact conditions for initiating common-envelope phases are not known (e.g. Ivanova et al., 2013; Pavlovskii et al., 2017; Temmink et al., 2023; Willcox et al., 2023), it is considered that low mass ratio systems are generally more unstable. This places a loose upper bound on the companion mass to be sufficiently smaller than the pre-common-envelope primary mass. On the other hand, the companion mass should be large enough such that the binary can stay bound after the SN. In the absence of NS natal kicks, the binary will survive when the ejecta mass is less than half of the total mass of the binary. The estimated ejecta mass for SN2022jli is $M_{\mathrm{ej}}\sim 1.5~M_{\odot}$ (Chen et al., 2024; Cartier et al., 2024), so assuming a NS mass of $M_{\mathrm{NS}}=1.4~M_{\odot}$, we cannot place a useful lower limit on the companion mass. Given these considerations, we choose a companion mass of $M_{2}=5~M_{\odot}$ as our fiducial model. For comparison, we also run models with companion masses of $M_{2}=3~M_{\odot}$ and $10~M_{\odot}$. All stellar models are constructed with the stellar evolution code MESA (v24.08.1; Paxton et al., 2011) with mostly default settings, assuming that the companion is close to its zero-age main-sequence structure.
For the orbital eccentricity, we choose $e=0.5$ as our baseline model and run $e=0.3,0.7$ and $0.8$ for comparison. All model parameters are summarized in Table 1 along with the periastron distance $a_{\mathrm{per}}$ and assumptions for the accretion and feedback (see Section 2.4). If we assume the pre-SN orbit was circular and that the ejecta mass was $M_{\mathrm{ej}}=1.5~M_{\odot}$, the mass loss alone can only excite an eccentricity of $e=M_{\mathrm{ej}}/(M_{\mathrm{NS}}+M_{2})\sim 0.34,0.23,0.13$ for $M_{2}=3,5,10~M_{\odot}$ respectively. Our choices of eccentricity implies that there was also a natal kick imparted to the NS with a magnitude of the order of $v_{\mathrm{kick}}\gtrsim 50~\mathrm{km~s}^{-1}$.
@@
Name $M_{2}$ $e$ $a_{\mathrm{per}}$ $p$ Feedback
@@
M05e05b 5 0.5 21.1 $0.0$ Bipolar
M05e03b 5 0.3 29.6 $0.0$ Bipolar
M05e07b 5 0.7 12.7 $0.0$ Bipolar
M10e05b 10 0.5 25.6 $0.0$ Bipolar
M03e05b 3 0.5 18.7 $0.0$ Bipolar
M05e05n 5 0.5 21.1 $0.0$ None
M05e05t 5 0.5 21.1 $0.0$ Thermal
M05e07b_r 5 0.7 12.7 $0.3$ Bipolar
M05e08b_r 5 0.8 8.44 $0.3$ Bipolar
@@ 2.2 Hydrodynamics code
We use a spherical coordinate system for all of our simulations. The companion star model is centered on the origin. Since our code cannot handle vacuum, we attach a dilute atmosphere outside the star, with a density profile $\rho=\rho_{0}(R_{2}/r)^{2}$, where $\rho_{0}$ is set to a value lower than the surface density of the companion. The outer boundary extends out to $r_{\mathrm{out}}=8\times 10^{15}~\mathrm{cm}\sim 115,000~R_{\odot}$, which is $\sim 3000$ times larger than the orbital semimajor axis. The radial grid spacing is chosen such that it is smallest around the surface of the star and increases in a geometric series inwards and outwards. This ensures that the pressure scale height in the star is always resolved by at least $\sim 10$ grid points [^3] Except for the very surface layers where the pressure scale height is prohibitively small.. We cover the star with $\sim 300$ radial grid points, and $N_{r}=800$ grid points for the entire domain. We assume equatorial symmetry, and thus only simulate the upper hemisphere where the polar grid spacing is uniform in $\cos\theta$, with $N_{\theta}=32$ grid points from pole to equator. This ensures that the solid angle per cell is uniform, minimizing asymmetries in numerical viscosity. It also places more resolution around the orbital plane, where most of the interesting dynamics are expected to take place. The azimuthal spacing is uniform with $N_{\phi}=256$ grid points. To relax the severe Courant conditions around the coordinate origin, we gradually reduce the angular resolution towards the centre, assuming full spherical symmetry ($N_{\theta}=N_{\phi}=1$) in the innermost 5 cells, $(N_{\theta},N_{\phi})=(2,16)$ in the next 5 cells, $(N_{\theta},N_{\phi})=(4,32)$ in the next 6 cells and so on.
@@ 2.3 SN-heating
After the SN explosion, the SN ejecta collide with the companion star, heating the surface layers through shocks. Only a small amount of mass is ejected in this process ($\lesssim 0.1~\%$) and the main effect is that the injected heat drives the surface out of thermal equilibrium, causing it to become overluminous and inflated (Hirai & Yamada, 2015; Hirai et al., 2018; Ogata et al., 2021; Chen et al., 2023; Hirai, 2023). Depending on the amount of energy injection and the stellar mass, the star can reach radii of a few $\times 100~R_{\odot}$. This is far beyond the orbital semi-major axis, inevitably causing direct interactions between the NS and the inflated envelope.
@@
Here, $R_{2}$ is the companion radius and $a$ is the orbital separation. The value of the energy deposition efficiency ranges around $p\sim 0.08$$0.12$, depending on how much the star gets compressed by the SN ejecta (see the discussion in Section 4.3 of Hirai et al. 2018). We use $p=1/12\sim 0.08$ for our simulations. Similarly, $m_{\mathrm{h}}$ is related to the ejecta mass $M_{\mathrm{ej}}$ through $m_{\mathrm{h}}=M_{\mathrm{ej}}\tilde{\Omega}/2$. The exact choice of $m_{h}$ does not influence the excess heat distribution as long as $m_{h}\ll M_{2}$.
@@ 2.4 Feedback method
The accretion and feedback is treated in our code as follows. For all cells within a specified accretion radius $|\bm{r}-\bm{r}_{\mathrm{NS}}|\leq R_{\mathrm{acc}}$, we assume that the mass is captured by the NS on the local dynamical timescale. The rate of mass and angular momentum captured by the NS is computed as
$\dot{M}_{\mathrm{cap}}$ $=\int_{|\bm{r}-\bm{r}_{\mathrm{NS}}|\leq R_{\mathrm{acc}}}\frac{\rho}{t_{\mathrm{ff}}}dV,$ (4)
$\dot{J}_{\mathrm{cap}}$ $=\int_{|\bm{r}-\bm{r}_{\mathrm{NS}}|\leq R_{\mathrm{acc}}}\frac{\rho}{t_{\mathrm{ff}}}(\bm{r}-\bm{r}_{\mathrm{NS}})\times(\bm{v}-\bm{v}_{\mathrm{NS}})dV,$ (5)
and the local free-fall time $t_{\mathrm{ff}}$ is defined as
@@
where $\phi_{\mathrm{soft}}(r)$ is the softened gravitational potential of the NS.
For most of our models, we optimistically assume that all of the captured material will be accreted onto the NS (optimistic model). However, based on magneto-hydrodynamic simulations of super-Eddington accretion, only a small fraction of the infalling material will actually accrete onto the compact object (e.g. Blandford & Begelman, 1999). Most of the material can be ejected via powerful disk winds before reaching the NS surface, not contributing much to the accretion luminosity. Therefore, assuming $\dot{M}_{\mathrm{NS}}=\dot{M}_{\mathrm{cap}}$ is likely too optimistic by orders of magnitude. As a more realistic choice, we run additional models where the mass accretion rate is scaled down as
$\dot{M}_{\mathrm{NS}}=\dot{M}_{\mathrm{cap}}\left(\frac{R_{\mathrm{NS}}}{R_{\mathrm{circ}}}\right)^{p},$ (7)
where $R_{\mathrm{circ}}\equiv\dot{J}_{\mathrm{cap}}^{2}/(GM_{\mathrm{NS}}\dot{M}_{\mathrm{cap}}^{2})$ is the circularization radius of the captured material (realistic model). Here, $G$ is the gravitational constant. For the value of $p$ we choose $p=0.3$, which is within the range of expected values (Blandford & Begelman, 1999; Inoue et al., 2023). The optimistic model is equivalent to $p=0$.
To represent accretion, we apply a source term to the continuity equation in the form
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\bm{v})=-\frac{\rho}{t_{\mathrm{ff}}}\left(\frac{R_{\mathrm{NS}}}{R_{\mathrm{circ}}}\right)^{p}.$ (8)
We then update the internal energy density of each cell such that the specific entropy is unchanged after applying the density source term. With the realistic mass accretion ($p=0.3$), we find that the accretion rate is typically more than an order of magnitude lower than that of the optimistic model ($p=0$). We choose higher orbital eccentricities in the realistic models than for the optimistic models (Table 1), which helps to compensate the luminosity decrease.
We apply several different methods to represent accretion feedback. We calculate the accretion luminosity as
@@
where we choose the NS radius to be $R_{\mathrm{NS}}=12~\mathrm{km}$.
In the thermal feedback model, we assume that all of the accretion energy is liberated in the form of radiation, and a fraction of it is captured by the gas and thermalizes. We define the optical depth within the accretion radius as
@@ 2.5 Light curve modelling
We use the post-processing module developed for HORMONE in Grishin et al. (2021) to compute synthetic light curves. It follows the ray-tracing method of Suzuki et al. (2016) where we solve the radiative transfer equations $dI_{\nu}/d\tau=-I_{\nu}+S_{\nu}$, where $I_{\nu},S_{\nu}$ are the frequency-dependent intensity and source function respectively, along rays pointing towards the observer. We assume $S_{\nu}=B_{\nu}(T)$, where $B_{\nu}$ is the Planck function for the local temperature and ignore any contributions of light scattered into the line of sight. The intensity is integrated over the whole projected area of the simulation domain and multiplied by $4\pi$ to obtain the isotropic-equivalent luminosity $L$. We separately carry out the same ray-tracing procedure for specific intensities at 6 different frequencies to get a rough spectrum. We fit the spectrum with a black body to compute an effective temperature $T_{\mathrm{eff}}$, which is then used to define a photospheric radius $R_{\mathrm{eff}}=\sqrt{L/(4\pi\sigma T_{\mathrm{eff}}^{4})}$ where $\sigma$ is the Stefan-Boltzmann constant.
@@
Given the high asymmetry of our problem, we also investigate the viewing direction dependence. Defining a Cartesian coordinate system with the orbital angular momentum vector as the $z$ axis, the direction opposite of the orbital eccentricity vector (Laplace-Runge-Lenz vector) as the $x$ axis, and the direction of the initial NS velocity as the $y$ axis, we choose 5 different viewing directions; +z,+x,x,+y,y+z,+x,-x,+y,-y (see the arrows in Figure 1). Note that the +z+z and z-z directions are identical as we assume equatorial symmetry. For the $\pm x,\pm y$ directions, we compute the light curve from two different polar angles, $\theta=\pi/2$ (edge on) and $\pi/4$ (off plane).
We do not simulate the SN explosion itself, which would be contributing significantly to the light curve, especially at early times. In reality the SN ejecta could also partially blanket the emission from the binary interactions we simulate, obscuring and/or smearing out short timescale features. To keep our results independent of the assumptions on the uncertain SN ejecta properties, we compute the light curves ignoring any effects of the SN ejecta. We discuss implications of this in Section 4.2.